### v1

# Friedrich-Hirzebruch-Allee 5

Abstract
Autonomous systems, such as self-driving cars, rely on reliable semantic environment perception for decision making. Despite great advances in video semantic segmentation, existing approaches ignore important inductive biases and lack structured and interpretable internal representations. In this work, we propose MCDS-VSS, a structured filter model that learns in a self-supervised manner to estimate scene geometry and ego-motion of the camera, while also estimating the motion of external objects. Our model leverages these representations to improve the temporal consistency of semantic segmentation without sacrificing segmentation accuracy. MCDS-VSS follows a prediction-fusion approach in which scene geometry and camera motion are first used to compensate for ego-motion, then residual flow is used to compensate motion of dynamic objects, and finally the predicted scene features are fused with the current features to obtain a temporally consistent scene segmentation. Our model parses automotive scenes into multiple decoupled interpretable representations such as scene geometry, ego-motion, and object motion. Quantitative evaluation shows that MCDS-VSS achieves superior temporal consistency on video sequences while retaining competitive segmentation performance. Animations and code to reproduce our results can be found in our project website.
1 Introduction
Video semantic segmentation (VSS) is the task of assigning a categorical label to each pixel in every frame of a video sequence [Zhou et al.(2022)Zhou, Porikli, Crandall, Van Gool, and Wang]. This task is highly relevant in the field of robotics, where understanding and interpreting scenes from video is crucial for many applications, e.g. autonomous driving [Siam et al.(2018)Siam, Gamal, Abdel-Razek, Yogamani, Jagersand, and Zhang] or indoor service tasks [Seichter et al.(2021)Seichter, Köhler, Lewandowski, Wengefeld, and Gross]. Thanks to the availability of high-quality image datasets, semantic segmentation of automotive scenarios has recently seen tremendous progress [Zhao et al.(2017)Zhao, Shi, Qi, Wang, and Jia, Chen et al.(2018)Chen, Zhu, Papandreou, Schroff, and Adam, Wang et al.(2020)Wang, Sun, Cheng, Jiang, Deng, Zhao, Liu, Mu, Tan, Wang, et al.]. However, obtaining temporally consistent segmentation of video sequences still remains a challenge due to the lack of large-scale annotated video datasets and the lack of suitable inductive biases for video processing.
To address these limitations, existing VSS models enforce temporal continuity by propagating features across multiple frames through the use of unstructured recurrent networks [Pfeuffer et al.(2019)Pfeuffer, Schulz, and Dietmayer, Siam et al.(2017)Siam, Valipour, Jagersand, and Ray], optical flow models [Ding et al.(2020)Ding, Wang, Zhou, Shi, Lu, and Luo, Gadde et al.(2017)Gadde, Jampani, and Gehler], or transformers [Wang et al.(2021a)Wang, Wang, and Liu, Li et al.(2021)Li, Wang, Chen, Niu, Si, Qian, and Zhang]; thus exploiting temporal correlations in the video sequences in a data-driven manner.
However, these models ignore specific properties from the target domain, which could potentially be incorporated into the model architecture in order to improve its performance and generalization capabilities. For instance, in the automotive domain, the observations taken from a moving vehicle can be decomposed into static background features, which move only due to the ego-motion of the vehicle, and dynamic object features that correspond to moving objects. Incorporating such motion and geometric inductive biases into the network architecture can lead to models producing a more temporally consistent interpretation of the scene, outperforming models that attempt to learn these properties solely from data.
To test this hypothesis, we propose MCDS-VSS, a structured recurrent model that explicitly incorporates geometry and motion inductive biases from the moving camera dynamic scene (MCDS) domain in order to improve the temporal consistency of a segmentation network. MCDS-VSS follows a prediction-fusion approach in which ego-motion is compensated by projecting scene features into the current time-step using estimated scene geometry and estimated camera motion. Estimated residual flow is then used to compensate for object motion. Finally, the predicted features are fused with the features extracted from the current frame to obtain a temporally consistent semantic segmentation of the scene.
Through self-supervised learning (SSL), MCDS-VSS learns to estimate scene geometry and ego-motion. It also estimates motion of additional objects (e.g. pedestrians or vehicles), and hard-wires our knowledge from the MCDS domain to project the previous scene features into the current time-step using these representations. The structured design of our filter allows us to factorize the perceived complex changes in the scene into simpler factors of variation; thus easing the modeling of temporal information.
Our experiments show that MCDS-VSS improves the temporal consistency of a segmentation model without compromising its segmentation performance, outperforming VSS baselines which ignore moving camera dynamic scene inductive biases, and performing comparably to state-of-the-art VSS models. Furthermore, MCDS-VSS parses an automotive scene into interpretable internal representations, such as depth, camera motion, and object flow.
In summary, our contributions are as follows:
  • We propose MCDS-VSS, a structured recurrent filter that improves the temporal consistency of a segmentation model without sacrificing segmentation performance.
  • MCDS-VSS learns depth and ego-motion in a self-supervised way, and uses these representations together with estimated object motion to propagate scene features.
  • Our model outperforms existing VSS baselines on Cityscapes—achieving superior temporal consistency and parsing the scene into human-interpretable representations.
2 Related Work
Video Semantic Segmentation:
VSS methods are often divided into two distinct categories. The first class aims to reduce the computational cost and improve the efficiency of segmentation models, instead of naively encoding and interpreting every single input frame. Several methods improve the efficiency by propagating and reusing features extracted from selected key frames [Jain et al.(2019)Jain, Wang, and Gonzalez, Zhu et al.(2017)Zhu, Xiong, Dai, Yuan, and Wei]; whereas other approaches achieve efficiency by employing lightweight neural network blocks [Pfeuffer and Dietmayer(2019), Shelhamer et al.(2016)Shelhamer, Rakelly, Hoffman, and Darrell] or by distilling the information from large teacher models to smaller models [Liu et al.(2020)Liu, Shen, Yu, and Wang]. The second category, to which our proposed method belongs, aims to improve the semantic segmentation performance and temporal consistency by exploiting the temporal continuity of video streams. Some methods exploit temporal dependencies between video frames and improve the consistency of the predicted segmentation maps by combining image segmentation models with recurrent neural networks (RNNs) [Siam et al.(2017)Siam, Valipour, Jagersand, and Ray, Pfeuffer et al.(2019)Pfeuffer, Schulz, and Dietmayer, Shelhamer et al.(2016)Shelhamer, Rakelly, Hoffman, and Darrell, Villar-Corrales et al.(2022)Villar-Corrales, Karapetyan, Boltres, and Behnke] or with attention-based modules [Wang et al.(2021a)Wang, Wang, and Liu, Lao et al.(2023)Lao, Hong, Guo, Zhang, Wang, Chen, and Chu, Sun et al.(2024)Sun, Liu, Ding, Wu, and Van Gool, Li et al.(2021)Li, Wang, Chen, Niu, Si, Qian, and Zhang, Sun et al.(2022)Sun, Liu, Tang, Chhatkuli, Zhang, and Van Gool]. Another family of works use an optical flow module to compute the feature correspondence between consecutive frames, and then use this flow for predictive feature learning [Ding et al.(2020)Ding, Wang, Zhou, Shi, Lu, and Luo, Zhu et al.(2017)Zhu, Xiong, Dai, Yuan, and Wei, Liu et al.(2020)Liu, Shen, Yu, and Wang, Xu et al.(2018)Xu, Fu, Yang, and Lee, Baghbaderani et al.(2024)Baghbaderani, Li, Wang, and Qi, Varghese et al.(2021)Varghese, Gujamagadi, Klingner, Kapoor, Bar, Schneider, Maag, Schlicht, Huger, and Fingscheidt, Hur and Roth(2016)].
Our method belongs to the latter category. However, unlike aforementioned approaches, MCDS-VSS incorporates assumptions from the domain of moving cameras and dynamic scenes into the model design, computing interpretable intermediate representations, which lead to more temporally consistent video segmentation results.
Improving Segmentation via Depth & Camera Motion Estimation:
Self-supervised depth estimation (SSDE) aims to learn the scene geometry from unlabeled monocular videos, without any recorded depth information. This is often achieved by training a neural network to jointly predict the scene depth and camera ego-motion between two video frames, synthesizing the second frame from the first using differentiable warping, and minimizing a photometric loss function [Garg et al.(2016)Garg, Bg, Carneiro, and Reid, Zhou et al.(2017)Zhou, Brown, Snavely, and Lowe, Bian et al.(2021)Bian, Zhan, Wang, Li, Zhang, Shen, Cheng, and Reid, Gordon et al.(2019)Gordon, Li, Jonschkowski, and Angelova, Wang et al.(2018)Wang, Buenaposada, Zhu, and Lucey].
The interplay between semantic segmentation and SSDE has been studied for various tasks, including depth estimation [Zama Ramirez et al.(2019)Zama Ramirez, Poggi, Tosi, Mattoccia, and Di Stefano], domain adaptation [Guizilini et al.(2021)Guizilini, Li, Ambrus, and Gaidon, Kuznietsov et al.(2022)Kuznietsov, Proesmans, and Van Gool, Wang et al.(2021b)Wang, Dai, Hoyer, Van Gool, and Fink], and semi-supervised learning [Hoyer et al.(2021)Hoyer, Dai, Chen, Koring, Saha, and Van Gool, Novosel et al.(2019)Novosel, Viswanath, and Arsenali]. These models exploit SSDE as an additional source of supervision, helping segmentation models learn high-level semantic features, especially when few labeled samples are available.
Several works have investigated the use of depth and motion for semantic segmentation in videos. Approaches like [Shi et al.(2020)Shi, Lin, Wang, Hung, and Wang, Athar et al.(2023)Athar, Li, Casas, and Urtasun, Cao and Behnke(2024), Zhu et al.(2023)Zhu, Han, Cai, Borse, Ghaffari, and Porikli] segment dynamic scenes by jointly processing video frames with depth information captured by LiDAR scanners; whereas methods like [Ma et al.(2017)Ma, Stückler, Kerl, and Cremers, Kundu et al.(2020)Kundu, Yin, Fathi, Ross, Brewington, Funkhouser, and Pantofaru, Chen et al.(2021)Chen, Lin, Wang, Yang, and Cheng, Rosu et al.(2020)Rosu, Quenzel, and Behnke] use depth and pose information to enforce consistency between segmentation maps from multiple viewpoints. Recently, depth-aware panoptic segmentation (PS) models [Qiao et al.(2021)Qiao, Zhu, Adam, Yuille, and Chen, Petrovai and Nedevschi(2023), Yuan et al.(2022)Yuan, Li, Yang, Cheng, Zhang, Tong, Zhang, and Tao] aim to jointly solve the tasks of PS and depth estimation by extending a segmentation model with a depth decoder and conditioning its prediction using instance-masks.
[…]
Figure 1: MCDS-VSS structured filter. Scene depth $\mathbf{d}_{t-1}$, ego-motion $\textrm{C}_{{t-1}}^{\,{t}}$, and object-motion $\textrm{F}_{t-1}^{\,t}$ are used to project scene features $\mathbf{s}_{t-1}$ to the current time $t$, where they are fused with current image features $\mathbf{h}_{t}$ to obtain a temporally consistent semantic segmentation $\mathbf{\hat{y}}_{t}$.
3 MCDS-VSS Structured Filtering Method
We propose MCDS-VSS, illustrated in Fig. 1, a structured filter that improves the temporal consistency of a semantic segmentation model on moving camera dynamic scene scenarios. MCDS-VSS learns in a self-supervised way to estimate geometry and motion representations, i.e., scene depth and camera ego-motion. It also estimates the motion of other agents in the scene, and uses these human-interpretable representations to propagate abstract scene features over time, thus improving its segmentation performance and temporal consistency.
MCDS-VSS is composed of an image encoder $\mathcal{E}_{\textrm{x}}$, a structured filter, and a segmentation decoder $\mathcal{D}_{\textrm{y}}$. It receives as input a sequence of RGB frames $\mathcal{X}=\{\mathbf{x}_{1},...,\mathbf{x}_{T}\}$ and encodes them into feature maps $\mathbf{h}_{1},...,\mathbf{h}_{T}\}$, which are then recursively processed to integrate temporal information, and decoded into semantic segmentation maps $\mathcal{Y}=\{\mathbf{\hat{y}}_{1}...,\mathbf{\hat{y}}_{T}\}$.
3.1 Learning of Geometry & Motion
MCDS-VSS learns in a self-supervised manner to estimate scene geometry and camera motion, which are then used to improve the temporal consistency of a segmentation model. Fig. 2 illustrates our two-step self-supervised approach for learning the scene geometry with camera motion and for distillation of object dynamics.
Scene Geometry and Ego-Motion: We train our model to predict the monocular depth and camera pose transformation of the vehicle in a self-supervised manner by solving a novel view-synthesis pretext task in which a target image $\mathbf{x}_{t}$ is rendered from a source image $\mathbf{x}_{t-1}$ by modeling the static scene features that change due to the ego-motion [Zhou et al.(2017)Zhou, Brown, Snavely, and Lowe, Godard et al.(2019)Godard, Mac Aodha, Firman, and Brostow].
To predict the scene geometry, MCDS-VSS incorporates a depth decoder $\mathcal{D}_{\textrm{d}}$ that outputs the depth $\mathbf{d}_{t}$ and inverse depth $\mathbf{d}^{\triangleleft}_{t}$111 With slight abuse of notation, we denote the $\mathbf{d}^{\triangleleft}_{t}$ as disparities. of the scene given the input feature maps $\mathbf{h}_{t}$; whereas to compute the camera motion between two images we employ a motion encoder $\mathcal{E}_{\textrm{m}}$ that computes motion features between two sets of feature maps, and an ego-motion decoder $\mathcal{D}_{\textrm{c}}$, which predicts the camera transformation between two time steps $\textrm{C}_{{t-1}}^{\,{t}}$, parameterized as a 6-dim. vector containing the translation and Euler angles of the camera transformation matrix. We then render the ego-warped image $\mathbf{\hat{x}}^{\textrm{ego}}_{t}$ using the estimated depth and ego-motion:
$\mathbf{\hat{x}}^{\textrm{ego}}_{t}=\mathcal{F}_{\textrm{fwd}}(\mathbf{x}_{t-1},\mathbf{d}_{t-1},\textrm{C}_{{t-1}}^{\,{t}},K),$ (1)
where $K\in{\mathbb{R}}^{3,3}$ are the camera intrinsic parameters, and $\mathcal{F}_{\textrm{fwd}}$ is the forward rendering function proposed in [Lee et al.(2021)Lee, Im, Lin, and Kweon]. If the depth and ego-motion estimates are accurate, the resulting warped images should match the target images except for occluded regions and moving objects. Therefore, as illustrated in Fig. 2(a), we train the modules for self-supervised depth and ego-motion estimation by optimizing the following loss function:
$\mathcal{L}_{\textrm{Depth}}=\mathcal{L}_{\textrm{Photo}}({\mathbf{\hat{x}}^{\textrm{ego}}_{t}},{\mathbf{x}_{t}})+\lambda_{\textrm{Reg}}\cdot\mathcal{L}_{\textrm{Reg}}({\mathbf{d}^{\triangleleft}_{t-1}}),$ (2)
$\mathcal{L}_{\textrm{Photo}}=\frac{\alpha}{2}(1-\text{SSIM}(\mathbf{\hat{x}}^{\textrm{ego}}_{t},\mathbf{x}_{t}))+(1-\alpha)||\mathbf{\hat{x}}^{\textrm{ego}}_{t}-\mathbf{x}_{t}||_{1},$ (3)
$\mathcal{L}_{\textrm{Reg}}=|\partial_{x}\widetilde{\mathbf{d}^{\triangleleft}_{t}}|e^{-|\partial_{x}\mathbf{x}_{t}|}+|\partial_{y}\widetilde{\mathbf{d}^{\triangleleft}_{t}}|e^{-|\partial_{y}\mathbf{x}_{t}|},$ (4)
where $\partial_{x}$ and $\partial_{y}$ are the spatial gradients in the $x$- and $y$-directions, SSIM is the structural similarity index, and $\widetilde{\mathbf{d}^{\triangleleft}}$ is the normalized disparity map. $\mathcal{L}_{\textrm{Photo}}$ is a photometric loss that measures the difference between the ego-warped and target images, and $\mathcal{L}_{\textrm{Reg}}$ is an edge-aware smoothing regularization [Godard et al.(2017)Godard, Mac Aodha, and Brostow] that encourages the normalized disparity maps to be locally smooth, except on the image edges. To mitigate the effect of disocclusions and moving objects during training, we use the auto-masking and per-pixel minimum processing steps proposed in [Godard et al.(2019)Godard, Mac Aodha, Firman, and Brostow].
[…]
(a) SSL of depth and ego-motion.
[…]
(b) Distillation of object-motion.
Figure 2: Learning geometry and motion. a) We learn the depth $\mathbf{d}_{t-1}$ and ego-motion $\textrm{C}_{{t-1}}^{\,{t}}$ given two video frames by enforcing a photometric loss between warped and target frame $\mathcal{L}_{\textrm{Photo}}$ and a depth regularization $\mathcal{L}_{\textrm{Reg}}$. b) Given an ego-warped image, we train a flow decoder to model the dynamics of moving objects by distilling a pretrained RAFT model.
Object Motion: Assuming static scenes as well as accurate depth and ego-motion estimates, the predicted ego-warped images $\mathbf{\hat{x}}^{\textrm{ego}}_{t}$ are identical, up to occluded regions, to the target images $\mathbf{x}_{t}$. Hence, we make the assumption that any major differences between such frames must be explained by external moving objects (e.g. driving cars or pedestrians).
As illustrated in Fig. 2(b), we estimate the residual optical flow $\hat{\textrm{F}}_{{t-1}}^{\,{t}}$ between the ego-warped and target images, which parameterizes the dynamics of moving objects, by training an residual flow decoder $\mathcal{R}_{\textrm{f}}$ while keeping all other modules frozen. $\mathcal{R}_{\textrm{f}}$ is trained to match the optical flow predictions of the large state-of-the-art optical flow model RAFT [Teed and Deng(2020)]:
$\mathcal{L}_{\textrm{Flow}}=||\hat{\textrm{F}}_{{t-1}}^{\,{t}}-\textrm{F}_{t-1}^{\,t}||_{1}.$ (5)
3.2 Structured Filter
The modules and representations described in
4 subsec: ssl
form the core of the MCDS-VSS structured filter, which is depicted in Fig. 1. It propagates information over time using two different filter states, namely a scene state $\mathbf{s}$ that encodes the scene contents and geometry and a camera state $\mathbf{c}$, which encodes the ego-motion of the vehicle.
It consists of six components: ego-motion filter, depth estimation, ego-motion compensation, residual flow estimation, object motion compensation, and feature fusion.
Ego-Motion Filter: The ego-motion filter extends the motion encoder and camera predictor modules in order to aggregate motion information over time and enforce the prediction of temporally consistent ego-motion. The temporal integration is achieved via a motion update module, which is implemented as a ConvGRU [Ballas et al.(2016)Ballas, Yao, Pal, and Courville] recurrent layer that jointly processes the motion features and previous camera state $\mathbf{c}_{t-1}$ and outputs an updated camera state $\mathbf{c}_{t}$ from which the ego-motion can be then predicted:
$\mathbf{c}_{t}=\text{ConvGRU}(\mathcal{E}_{\textrm{m}}(\mathbf{h}_{t},\mathbf{h}_{t-1}),\mathbf{c}_{t-1}),\hskip 28.45274pt\textrm{C}_{{t-1}}^{\,{t}}=\mathcal{D}_{\textrm{c}}(\mathbf{c}_{t}).$ (6)
Depth Estimation and Ego-Motion Compensation: Given the past scene state $\mathbf{s}_{t-1}$, the depth decoder $\mathcal{D}_{\textrm{d}}$ computes the depth map $\mathbf{d}_{t-1}$. This scene geometry and the ego-motion $\textrm{C}_{{t-1}}^{\,{t}}$ are used as in Eq. (1) to project $\mathbf{s}_{t-1}$ to time $t$. The resulting ego-warped scene state $\mathbf{s}^{\textrm{ego}}_{t}$ encodes scene contents and geometry after compensation for ego-motion.
Object-Motion Compensation: This module models dynamic objects in the scene, such as pedestrians or vehicles, and updates the scene state to compensate for the motion of such objects. We jointly process the ego-warped scene state $\mathbf{s}^{\textrm{ego}}_{t}$ and the current image features $\mathbf{h}_{t}$ with the residual flow decoder $\mathcal{R}_{\textrm{f}}$ in order to compute the residual flow $\hat{\textrm{F}}_{{t-1}}^{\,{t}}$, which represents the pixel displacement of moving objects between consecutive time steps. The ego-warped features $\mathbf{s}^{\textrm{ego}}_{t}$ are then projected into the current time-step by applying the displacement encoded in the residual flow map, followed by bilinear interpolation to obtain valid coordinate values. The resulting features $\mathbf{s}^{\textrm{full}}_{t}$ not only incorporate the motion of dynamic objects in the scene, but can also correct alignment errors between $\mathbf{s}^{\textrm{ego}}_{t}$ and $\mathbf{h}_{t}$ that might occur due to inaccurate depth or ego-motion estimates.
Feature Fusion: While the previous modules propagate scene features over time, the feature fusion module allows MCDS-VSS to combine the projected scene features $\mathbf{s}^{\textrm{full}}_{t}$ with the observed encoded features $\mathbf{h}_{t}$. This fusion operation is performed by an update gate mask $\mathbf{u}\in[0,1]$, which determines in a data-driven manner for each feature map and spatial location whether one can rely on the current features $\mathbf{h}_{t}$ or on prior knowledge propagated through the filter $\mathbf{s}^{\textrm{full}}_{t}$. The resulting fused scene features $\mathbf{s}_{t}$ are part of the next state of the MCDS-VSS filter. More formally, this process can be described as:
$\mathbf{u}=\sigma(\text{Conv}_{\textrm{s}}(\mathbf{s}^{\textrm{full}}_{t})+\text{Conv}_{\textrm{h}}(\mathbf{h}_{t})+b,)$ (7)
$\mathbf{s}_{t}=\mathbf{u}\odot\mathbf{s}^{\textrm{full}}_{t}+(1-\mathbf{u})\odot\mathbf{h}_{t},$ (8)
where $\text{Conv}_{\textrm{s}}$ and $\text{Conv}_{\textrm{h}}$ are convolution blocks, $b$ a learned bias, and $\sigma$ the sigmoid function.
4.1 Model Training
MCDS-VSS consists of multiple components addressing different subtasks: depth estimation, ego-motion estimation, ego-motion compensation, object motion estimation, object motion compensation, and feature fusion. Naively training such a model in an end-to-end manner with a video segmentation objective can result in bad local optima, where the model does not learn interpretable representations (e.g. depth or object flow).
To ease the training process, we propose a multi-stage approach in which we first train the encoder and decoder modules using image pairs or triplets, and then integrate and train the filter modules using sequences of six frames in order to improve segmentation performance and temporal consistency while retaining interpretable representations.
Table 1: Training stages and hyper-parameters.
Stage Training Goal Loss Function LR # Imgs
1 Segmentation & SSL Geometry $\mathcal{L}_{\textrm{Seg}}+\lambda_{\textrm{D}}\cdot\mathcal{L}_{\textrm{Depth}}$   (2) $2\cdot 10^{-4}$ 3
2 Distillation of Object Motion $\mathcal{L}_{\textrm{Flow}}$   (5) $1\cdot 10^{-4}$ 2
3 Ego-Motion Filter $\mathcal{L}_{\textrm{Ego}}$   (9) $8\cdot 10^{-5}$ 6
4 Temporal Integration $\mathcal{L}_{\textrm{Seg}}+\lambda_{\textrm{TC}}\cdot\mathcal{L}_{\textrm{TC}}$   (10) $8\cdot 10^{-5}$ 6
MCDS-VSS undergoes a four-stage training process, outlined in Table 1. Initially, as detailed in
5 subsec: ssl
, MCDS-VSS encoder and decoder modules are jointly trained for self-supervised learning of geometry and ego-motion, as well as for semantic segmentation by minimizing a combination of cross entropy $\mathcal{L}_{\textrm{Seg}}$ and SSL geometry $\mathcal{L}_{\textrm{Depth}}$ losses. Following [Godard et al.(2019)Godard, Mac Aodha, Firman, and Brostow], we use image triplets $(\mathbf{x}_{t-\tau},\mathbf{x}_{t},\mathbf{x}_{t+\tau})$, with $\mathbf{x}_{t}$ being the target. Subsequently, we train the residual flow decoder as described in
6 subsec: ssl
while keeping the remaining modules frozen. In the third stage, with the goal of improving the temporal continuity of the predicted ego-motion, we train the ego-motion filter and compensation by minimizing the loss function:
$\mathcal{L}_{\textrm{Ego}}=\frac{1}{T}\sum_{t=1}^{T}\mathcal{L}_{\textrm{Photo}}({\mathbf{\hat{x}}^{\textrm{ego}}_{t}},{\mathbf{x}_{t}}),$ (9)
which enforces the model to compute accurate camera motion estimates in order to align the ego-warped state with the current observations. Finally, we jointly train the fusion module and fine-tune the segmentation decoder by minimizing the following loss function:
$\mathcal{L}=\frac{1}{T}\sum_{t=1}^{T}\mathcal{L}_{\textrm{Seg}}({\mathbf{\hat{y}}_{t}},{\mathbf{y}_{t}})+\lambda_{\textrm{TC}}\cdot\mathcal{L}_{\textrm{TC}}({\mathbf{\tilde{y}}_{t}},{\mathbf{\hat{y}}_{t}}),$ (10)
where $\mathcal{L}_{\textrm{Seg}}$ is the cross entropy loss function and $\mathcal{L}_{\textrm{TC}}$ is a temporal consistency regularizer that enforces the segmentation $\mathbf{\tilde{y}}_{t}$ computed by decoding $\mathbf{s}^{\textrm{full}}_{t}$ to be close to the actual predicted segmentation maps $\mathbf{\hat{y}}_{t}$.
7 Experimental Evaluation
7.1 Experiment Setup
Dataset: We evaluate MCDS-VSS on the Cityscapes [Cordts et al.(2016)Cordts, Omran, Ramos, Rehfeld, Enzweiler, Benenson, Franke, Roth, and Schiele] dataset, which contains 5,000 automotive video sequences, each containing 30 images of size 1024$\times$2048, recorded in 50 German cities, where only the 20th frame is annotated. This dataset is a good benchmark for our model, since it contains real-world dynamic scenes recorded from a moving vehicle. We augment the data using color jittering, mirroring, and random cropping.
Evaluation Metrics: We evaluate the segmentation performance and temporal consistency of our model. The performance is evaluated using the mean Intersection-over-Union (mIoU). Following [Liu et al.(2020)Liu, Shen, Yu, and Wang], we measure the temporal consistency (TC) of our predicted segmentation maps by computing the mean flow warping error between every two neighboring frames. Our results are computed using single-scale testing on the full image resolution.
Implementation Details: We train two MCDS-VSS variants using distinct image encoder and segmentation decoder architectures. Namely, a small variant based on DeepLabV3+ [Chen et al.(2018)Chen, Zhu, Papandreou, Schroff, and Adam] with ResNet18 [He et al.(2016)He, Zhang, Ren, and Sun] backbone, and a larger variant based on HRNetV2 [Wang et al.(2020)Wang, Sun, Cheng, Jiang, Deng, Zhao, Liu, Mu, Tan, Wang, et al.]. The depth and pose decoders closely follow [Godard et al.(2019)Godard, Mac Aodha, Firman, and Brostow], which output inverse depth maps and a 6-dimensional vector containing the camera translation and Euler angles, respectively. Finally, our residual flow decoder is a lightweight version of RAFT [Teed and Deng(2020)], for which, to integrate into our filter, we replace the context and feature encoders with a single convolutional block. We emphasize that MCDS-VSS is architecture-agnostic and could be implemented with different model designs. Further implementation details are provided in Appendix B.
7.2 Comparison with Existing Methods
In Table 2, we quantitatively compare MCDS-VSS with several existing image and video segmentation models using small (left) and larger (right) backbones. For both variants, MCDS-VSS achieves the highest temporal consistency among all compared methods, while retaining a competitive segmentation performance. Furthermore, in contrast to other approaches aiming to improve the TC of a segmentation model, e.g. TCNet [Varghese et al.(2021)Varghese, Gujamagadi, Klingner, Kapoor, Bar, Schneider, Maag, Schlicht, Huger, and Fingscheidt], MCDS-VSS does not sacrifice segmentation performance in order to improve the temporal consistency, outperforming multiple VSS models for both backbone variants.
Table 2: Comparison of segmentation models on the Cityscapes validation set using different backbones. Best two results are highlighted in boldface and underlined, respectively.
Cityscapes
[…]
DeepLabV3+ [Chen et al.(2018)Chen, Zhu, Papandreou, Schroff, and Adam] ResNet18 75.2 69.8
Accel [Jain et al.(2019)Jain, Wang, and Gonzalez] ResNet18 72.1 70.3
SKD [Liu et al.(2019)Liu, Chen, Liu, Qin, Luo, and Wang] ResNet18 74.5 68.2
ETC [Liu et al.(2020)Liu, Shen, Yu, and Wang] PSPNet18 73.1 70.6
ETC [Liu et al.(2020)Liu, Shen, Yu, and Wang] MobileNetV2 73.9 69.9
TCNet [Varghese et al.(2021)Varghese, Gujamagadi, Klingner, Kapoor, Bar, Schneider, Maag, Schlicht, Huger, and Fingscheidt] ResNet18 62.2 72.1
TDNet [Hu et al.(2020)Hu, Caba, Wang, Lin, Sclaroff, and Perazzi] BiSeNet18 75.0 70.2
TDNet [Hu et al.(2020)Hu, Caba, Wang, Lin, Sclaroff, and Perazzi] PSPNet18 76.8 70.4
STT [Li et al.(2021)Li, Wang, Chen, Niu, Si, Qian, and Zhang] BiSeNet18 75.8 71.4
STT [Li et al.(2021)Li, Wang, Chen, Niu, Si, Qian, and Zhang] ResNet18 77.3 73.0
MCDS-VSS (ours) ResNet18 75.1 74.5
[…]
Figure 3: Qualitative evaluation. a) Input frames, b) HRNetV2, c) MCDS-VSS (ours), d) Estimated scene depth, e) Estimated residual flow. We highlight areas of the segmentation masks where MCDS-VSS obtains visibly more accurate and consistent segmentations.
In Fig. 3, we show a qualitative result comparing MCDS-VSS with an HRNetV2 baseline. Our method achieves more accurate and temporally consistent segmentations compared to HRNet, correctly segmenting the traffic signs and reducing the amount of flickering between frames. Furthermore, we show MCDS-VSS interpretable intermediate representations, such as the depth of the scene and the residual flow, which encodes the movement of the pedestrians in the scene, as well as some corrections for other objects, such as the hood of the ego-vehicle. Further visualizations can be found in Appendix D.
7.3 Ablation Study
To understand the effectiveness of MCDS-VSS, we ablate our filter design and measure the contribution of different steps in our training process.
Filter Design: Given the same DeepLabV3+ model trained with the SSL procedure described in
8 subsec: ssl
, we compare MCDS-VSS with different filter designs, including unstructured RNNs (ConvGRU) [Siam et al.(2017)Siam, Valipour, Jagersand, and Ray, Pfeuffer et al.(2019)Pfeuffer, Schulz, and Dietmayer], optical-flow based filters [Gadde et al.(2017)Gadde, Jampani, and Gehler] (flow-only), and a MCDS-VSS variant modeling only the static scene features (geom-only) [Wagner et al.(2018)Wagner, Fischer, Herman, and Behnke]. The results, reported in Table 3, show that MCDS-VSS, which decouples the modeling of static and dynamic scene features, achieves the best segmentation performance and temporal consistency.
Model Ablation: In Table 4 we measure the effect that our SSL training and MCDS-VSS filter have on the segmentation performance and temporal consistency. For two different segmentation models, i.e. DeepLabV3+ and HRNetV2, we compare the results after each training stage with those of the model trained for image segmentation only. First, we note that SSL training improves the temporal consistency without significantly compromising the segmentation performance. We argue that jointly learning SSL depth and motion with semantic segmentation allows the model to encode the input frames into more robust geometry-aware representations. Finally, the MCDS-VSS filter significantly improves the temporal consistency (>4.6% w.r.t base model), while almost matching the segmentation performance of the base DeepLabV3+ model, and even outperforming HRNetV2.
Table 3: Comparison of various filter designs. We highlight the diff. to baseline.
Results
[…]
ResNet18 + SSL 74.76 70.73
  + ConvGRU [Siam et al.(2017)Siam, Valipour, Jagersand, and Ray] 73.37 (-1.39) 69.64 (-1.09)
  + Flow-Only [Gadde et al.(2017)Gadde, Jampani, and Gehler] 74.95 (+0.19) 73.19 (+2.46)
  + Geom-Only [Wagner et al.(2018)Wagner, Fischer, Herman, and Behnke] 74.90 (+0.14) 73.80 (+3.07)
  + MCDS-VSS 75.07 (+0.31) 74.53 (+3.80)
Table 4: Effect of SSL geometry & motion and MCDS-VSS. We highlight the diff. to baseline.
Results
[…]
ResNet18 75.17 69.89
  + SSL Geom. & Motion 74.76 (-0.44) 70.73 (+0.91)
  + MCDS-VSS 75.07 (-0.10) 74.53 (+4.64)
HRNetV2 76.32 70.58
  + SSL Geom. & Motion 76.45 (+0.13) 71.65 (+1.07)
  + MCDS-VSS 77.14 (+0.82) 75.34 (+4.76)
9 Conclusion
We proposed MCDS-VSS, a structured recurrent model for VSS, which learns in a self-supervised manner to estimate scene geometry and camera ego-motion. It also estimates the motion of external objects and leverages these representations to improve the temporal consistency of a semantic segmentation model without sacrificing segmentation performance. MCDS-VSS follows a prediction-fusion approach in which scene geometry and camera motion are first used to compensate for ego-motion, then residual flow is used to compensate the motion of dynamic objects, and finally the projected features are fused with the current observations. In our experiments, we showed that MCDS-VSS outperforms multiple VSS baselines on Cityscapes—achieving superior temporal consistency and parsing the scene into human-interpretable representations.
Acknowledgments
This work was funded by grant BE 2556/16-2 (Research Unit FOR 2535 Anticipating Human Behavior) of the German Research Foundation (DFG).
Appendix A Evaluation Metrics
To evaluate MCDS-VSS, we compute its segmentation performance, temporal consistency, throughput and inference speed.
Following the standard practice, we use the mean Intersection over Union (mIoU) to evaluate the segmentation performance.
[…]
We measure the throughput and inference speed of our model in frames per second (FPS) and milliseconds (ms), respectively. For this purpose, we perform inference with MCDS-VSS on 200 different video sequences of 6 frames and average the throughput and time across frames and sequences.
Appendix B Implementation Details
B.1 Network Details
In this section, we describe the network architectures and operation of each module in MCDS-VSS. We emphasize that MCDS-VSS is architecture-agnostic and could be implemented with different, e.g. more powerful or efficient, model designs.
Image Encoder & Segmentation Decoder: We implement two distinct MCDS-VSS variants, whose image encoder and segmentation decoder follow the architecture of two popular image semantic segmentation models, namely DeepLabV3+ [Chen et al.(2018)Chen, Zhu, Papandreou, Schroff, and Adam] with a ResNet18 [He et al.(2016)He, Zhang, Ren, and Sun] backbone and HRNetV2 [Wang et al.(2020)Wang, Sun, Cheng, Jiang, Deng, Zhao, Liu, Mu, Tan, Wang, et al.] with a channel multiplier of 18, respectively. In both cases we initialize the parameters of the encoders with those of the model pretrained on ImageNet, whereas the segmentation decoders are initialized with random weights.
Motion Encoder: The motion encoder concatenates the image features from two consecutive time steps (i.e. $\mathbf{h}_{t-1}$ and $\mathbf{h}_{t}$) across the channel dimension, and processes this representation with three convolutional layers, followed by batch normalization and ReLU activation functions.
Table 5: Network architecture of the depth decoder module.
Layer Modules Output Dim.
Block 1 Conv + ELU 256 $\times$ H/8 $\times$ W/8
Conv + ELU 256 $\times$ H/8 $\times$ W/8
Block 2 Conv + ELU 128 $\times$ H/8 $\times$ W/8
Conv + ELU 128 $\times$ H/8 $\times$ W/8
Block 3 Conv + ELU + Ups. 64 $\times$ H/4 $\times$ W/4
Conv + ELU 64 $\times$ H/4 $\times$ W/4
Block 4 Conv + ELU + Ups. 32 $\times$ H/2 $\times$ W/2
Conv + ELU 32 $\times$ H/2 $\times$ W/2
Block 5 Conv + ELU + Ups. 16 $\times$ H $\times$ W
Conv + ELU 16 $\times$ H $\times$ W
Disp. Pred. Conv + Sigmoid 1 $\times$ H $\times$ W
Table 6: Network architecture of the ego-motion decoder module.
Layer Modules Output Dim.
Block 1 Conv + BN + ReLU 256 $\times$ H/8 $\times$ W/8
Block 2 Conv + BN + ReLU 256 $\times$ H/8 $\times$ W/8
Block 3 Conv + BN + ReLU 128 $\times$ H/8 $\times$ W/8
Block 4 Conv + Pool + ReLU 128 $\times$ H/16 $\times$ W/16
Ego Motion Conv 6 $\times$ H/16 $\times$ W/16
Global Avg. Pool 6
Depth Decoder: The network architecture of our depth decoder, which is reported in Table 5, closely follows [Godard et al.(2017)Godard, Mac Aodha, and Brostow, Godard et al.(2019)Godard, Mac Aodha, Firman, and Brostow]. It is composed of five convolutional blocks using reflection padding, followed by ELU nonlinearities. Each of the last three convolutional blocks upsamples the feature maps by a factor of two using nearest-neighbor upsampling. The depth decoder outputs normalized inverse depth maps $\mathbf{d}^{\triangleleft}$, which are then converted into depth maps $\mathbf{d}$ by:
$\frac{1}{\mathbf{d}}=\frac{1}{D_{\mathrm{min}}}+(\frac{1}{D_{\mathrm{max}}}-\frac{1}{D_{\mathrm{min}}})\cdot\mathbf{d}^{\triangleleft}\,,$ (11)
where $D_{\mathrm{min}}$ and $D_{\mathrm{max}}$ are constant values defining the minimum and maximum depth values in the scene, set to $D_{\mathrm{min}}=0.1$m and $D_{\mathrm{max}}=100$m for the Cityscapes dataset.
Ego-Motion Decoder: The ego-motion decoder, which is reported in Table 6, processes the motion features with a series of convolution layers, followed by batch normalization and ReLU nonlinearities. The final layer outputs a 6-dimensional vector representing the translation and rotation (parameterized as Euler angles) of the camera transformation matrix.
Residual Flow Decoder: The residual flow decoder is implemented as a modified lightweight version of RAFT [Teed and Deng(2020)]. To seamlessly integrate this module into our MCDS-VSS filter, we modify the implementation of raft_small provided by PyTorch [^2] https://pytorch.org/vision/main/models/raft.html by replacing the expensive feature and context encoders with a single convolutional block that directly processes the ego-warped $\mathbf{s}^{\textrm{ego}}_{t}$ and image features $\mathbf{h}_{t}$. We set the number of refinement iterations to 10.
MCDS-VSS Filter: We instantiate the MCDS-VSS filter using the modules described above, after being pretrained for semantic segmentation, SSL of depth and ego-motion, and distillation of object motion. For the first image in a video sequence, MCDS-VSS directly predicts its semantic segmentation, without the use of any temporal filtering. For all other frames, MCDS-VSS employs the structured filtering method described in the paper. The scene feature state is initialized with the image features from the first frame in the video sequence ($\mathbf{s}_{1}=\mathbf{h}_{1}$), whereas the initial camera state $\mathbf{c}_{1}$ is initialized with zeros. We experimented with learning the initial state representations; however it did not yield any qualitative or quantitative improvements, while increasing the number of learnable parameters.
Table 7: Training stages and hyper-parameters.
Stage Training Goal Loss Function LR Batch Train Time # Imgs
1 Segmentation & SSL Geometry $\mathcal{L}_{\textrm{Seg}}+\lambda_{\textrm{D}}\cdot\mathcal{L}_{\textrm{Depth}}$ $2\cdot 10^{-4}$ 12 40h 3
2 Distillation of Object Motion $\mathcal{L}_{\textrm{Flow}}$ $1\cdot 10^{-4}$ 4 18h 2
3 Ego-Motion Filter $\mathcal{L}_{\textrm{Ego}}$ $8\cdot 10^{-5}$ 8 8h 6
4 Temporal Integration $\mathcal{L}_{\textrm{Seg}}+\lambda_{\textrm{TC}}\cdot\mathcal{L}_{\textrm{TC}}$ $8\cdot 10^{-5}$ 4 12h 6
B.2 Training and Inference
All our models are implemented in PyTorch [Paszke et al.(2017)Paszke, Gross, Chintala, Chanan, Yang, DeVito, Lin, Desmaison, Antiga, and Lerer] and trained with two NVIDIA A100 (80GB) GPUs. For each of the four training stages undergone by MCDS-VSS, we report in Table 7 the most relevant hyper-parameters, including the approximate training time, learning rate, batch size and number of images per sequence. We empirically set the loss weight values to $\lambda_{\textrm{D}}=2$, $\lambda_{\textrm{Reg}}=10^{-5}$ and $\lambda_{\textrm{TC}}=1$.
Tables 8 and 9 report the number of learnable parameters, throughput and inference time for each individual module, as well as for the complete model, for the MCDS-VSS variants based on DeepLabV3+ and HRNetV2, respectively. We emphasize that the ego-motion and object motion compensation modules do not have any learnable parameters, but instead hard-wire our knowledge from the moving camera dynamic scene domain to project the previous scene features into the current time-step using geometry and motion representations. We also observe that MCDS-VSS inference is severely limited by its residual flow decoder. Adapting such module to exploit recent advances in fast optical flow estimation [Kong et al.(2021)Kong, Shen, and Yang]), as well as using more efficient image encoders [Howard et al.(2019)Howard, Sandler, Chu, Chen, Chen, Tan, Wang, Zhu, Pang, Vasudevan, et al.], could allow MCDS-VSS to be used for real time video semantic segmentation.
Table 8: Throughput, inference speed (in ms) and number of learnable parameters for MCDS-VSS based on DeepLabV3+.
Model # Params. FPS Inf. (ms)
Image Enc   $\mathcal{E}_{\textrm{x}}$ 15.3M 76.9 12.9
Motion Enc   $\mathcal{E}_{\textrm{m}}$ 4.8M 279.4 3.6
Motion Update 2.8M 476.2 2.1
Ego-Motion Dec   $\mathcal{D}_{\textrm{c}}$ 1.6M 492.8 2.0
Depth Dec   $\mathcal{D}_{\textrm{d}}$ 1.9M 227.1 4.4
Ego-Motion Comp. 0 63.4 15.8
Residual Flow Dec   $\mathcal{R}_{\textrm{f}}$ 2.7M 14.0 71.8
Object Motion Comp. 0 775.6 1.3
Feature Fusion 2.6M 215.3 4.6
Segmentation Dec  $\mathcal{D}_{\textrm{y}}$ 1.3M 196.7 5.1
Total MCDS-VSS 32.9M 9.0 111.6
Table 9: Throughput, inference speed (in ms) and number of learnable parameters for MCDS-VSS based on HRNetV2.
Model # Params. FPS Inf. (ms)
Image Enc   $\mathcal{E}_{\textrm{x}}$ 9.5M 36.7 27.2
Motion Enc   $\mathcal{E}_{\textrm{m}}$ 5.0M 75.8 13.2
Motion Update 2.8M 166.7 6.0
Ego-Motion Dec   $\mathcal{D}_{\textrm{c}}$ 1.6M 174.4 5.7
Depth Dec   $\mathcal{D}_{\textrm{d}}$ 1.9M 62.4 16.0
Ego-Motion Comp. 0 47.9 20.9
Residual Flow Dec   $\mathcal{R}_{\textrm{f}}$ 2.9M 13.4 74.6
Object Motion Comp. 0 886.2 1.1
Feature Fusion 2.6M 76.8 13.0
Segmentation Dec   $\mathcal{D}_{\textrm{y}}$ 78.8K 627.6 1.6
Total MCDS-VSS 26.2M 5.6 177.6
Appendix C Additional Quantitative Results
In Table 10 we compare for individual classes of the Cityscapes dataset the segmentation performance and temporal consistency of MCDS-VSS with an HRNetV2 baseline trained for semantic segmentation and SSL of geometry and motion. MCDS-VSS achieves the best segmentation performance and temporal consistency for most classes in the dataset, especially for those corresponding to moving objects, such as car, truck, bus or train.
Table 10: Segmentation performance (mIoU) and temporal consistency (TC) for individual Cityscapes classes. We compare MCDS-VSS with HRNetV2 backbone with an HRNetV2 model trained for semantic segmentation and SSL of geometry and motion.
Road Sidewalk Building Wall Fence Pole Traf. Light Traf. Sign Person Rider Car Truck Bus Train Motorbike Bicycle Mean
mIoU Baseline 98.1 84.3 92.5 50.3 59.2 66.5 70.3 79.3 81.8 63.5 94.6 72.4 83.5 67.7 61.4 76.6 76.5
Ours 98.1 84.7 92.7 51.3 59.6 65.9 70.5 79.3 81.6 63.7 94.9 77.6 85.0 70.6 61.7 76.7 77.1
TC Baseline 98.7 86.0 91.1 60.8 60.1 67.1 62.4 77.1 69.0 65.8 89.0 58.1 66.8 45.8 49.4 66.6 71.7
Ours 98.9 88.2 92.1 64.1 65.9 69.1 69.2 79.5 70.5 67.8 90.5 65.2 69.6 48.3 53.6 71.0 75.3
Appendix D Additional Qualitative Results
D.1 Effect of Each MCDS-VSS Stage
In Fig. 4 we display the semantic segmentations obtained when decoding the scene features from different stages of our MCDS-VSS filter. We can observe how the segmentations after ego-motion compensation (Fig. 4 b) atone for the movement of the ego-vehicle, correctly representing static scene features such as buildings or the bicycle. However, the dynamics of moving objects (e.g. yellow car) are not addressed in this step, thus not compensating for such movement. This limitation is addressed in the object motion compensation step (Fig. 4 c). However, the disocclusions resulting from the moving car and inaccuracies in the residual flow estimation can lead to segmentation errors. Finally, fusing the projected scene state features with the current observations (Fig. 4 d) leads to a more accurate segmentation of the scene. In (Fig. 4 e) we visualize the update gate mask $\mathbf{u}$ employed for feature fusion. For visualization purposes, we take the mean across all channels and assign lighter colors to spatial locations where MCDS-VSS relies on the propagated state $\mathbf{s}^{\textrm{full}}_{t}$, whereas darker colors correspond to the current features $\mathbf{h}_{t}$. We observe that MCDS-VSS relies on the scene state to represent static areas such as buildings or the street, whereas it relies heavier on observations for accurately segmenting fast moving objects or thin structures (e.g. poles or street signs).
[…]
Figure 4: Video segmentation for each stage in MCDS-VSS. a) Input images, b) segmentation after ego-motion compensation, c) segmentation after object motion compensation, d) segmentation after feature fusion, e) feature fusion update mask, lighter colors mean that filter information is used, whereas darker ones correspond to observations.
[…]
Figure 5: MCDS-VSS qualitative evaluation. a) Input frames, b) semantic segmentation, c) scene depth, d) residual flow.
[…]
Figure 6: MCDS-VSS qualitative evaluation. a) Input frames, b) semantic segmentation, c) scene depth, d) residual flow.
[…]
Figure 7: MCDS-VSS qualitative evaluation. a) Input frames, b) semantic segmentation, c) scene depth, d) residual flow.
[…]
Figure 8: MCDS-VSS qualitative evaluation. a) Input frames, b) semantic segmentation, c) scene depth, d) residual flow.
D.2 MCDS-VSS Qualitative Results
Despite being trained without any depth or motion labels, and solely with semantic segmentation annotations for a single frame per video, our model parses the scene into semantic, geometry and motion representations. In Figs. 58, we show on four validation sequences how MCDS-VSS obtains an accurate and temporally consistent segmentation of the scene, estimates the scene depth, and computes the residual motion flow, which encodes the movement of dynamic objects in the scene, as well as some minor motion corrections for other objects and scene features.
D.3 Point Clouds
Figs. 912 show examples of RGB and semantic point clouds rendered by backprojecting image values and semantic labels using the depth maps estimated by MCDS-VSS and known camera intrinsics. The high-quality depth maps computed by MCDS-VSS allow for an accurate 3D representation of the scene.
[…]
Figure 9: RGB and semantic point clouds rendered by lifting image values and semantic labels to 3D space.
[…]
Figure 10: RGB and semantic point clouds rendered by lifting image values and semantic labels to 3D space.
[…]
Figure 11: RGB and semantic point clouds rendered by lifting image values and semantic labels to 3D space.
[…]
Figure 12: RGB and semantic point clouds rendered by lifting image values and semantic labels to 3D space.

### v2

@@
-Friedrich-Hirzebruch-Allee 5
+Center for Robotics and
 Abstract
@@
 Quantitative evaluation shows that MCDS-VSS achieves superior temporal consistency on video sequences while retaining competitive segmentation performance.
-Animations and code to reproduce our results can be found in our project website.
+Code and pretrained models are available in the project website.
 1 Introduction
@@
 Through self-supervised learning (SSL), MCDS-VSS learns to estimate scene geometry and ego-motion.
-It also estimates motion of additional objects (e.g.
+It also estimates motion of additional moving objects (e.g.
 pedestrians or vehicles), and hard-wires our knowledge from the MCDS domain to project the previous scene features into the current time-step using these representations.
@@
 The first class aims to reduce the computational cost and improve the efficiency of segmentation models, instead of naively encoding and interpreting every single input frame.
-Several methods improve the efficiency by propagating and reusing features extracted from selected key frames [Jain et al.(2019)Jain, Wang, and Gonzalez, Zhu et al.(2017)Zhu, Xiong, Dai, Yuan, and Wei]; whereas other approaches achieve efficiency by employing lightweight neural network blocks [Pfeuffer and Dietmayer(2019), Shelhamer et al.(2016)Shelhamer, Rakelly, Hoffman, and Darrell] or by distilling the information from large teacher models to smaller models [Liu et al.(2020)Liu, Shen, Yu, and Wang].
+Several methods improve the efficiency by propagating and reusing features extracted from selected key frames [Jain et al.(2019)Jain, Wang, and Gonzalez, Zhu et al.(2017)Zhu, Xiong, Dai, Yuan, and Wei]; whereas other approaches achieve efficiency by employing lightweight neural network blocks [Pfeuffer and Dietmayer(2019), Shelhamer et al.(2016)Shelhamer, Rakelly, Hoffman, and Darrell] or by distilling the information from large teacher models into smaller models [Liu et al.(2020)Liu, Shen, Yu, and Wang].
 The second category, to which our proposed method belongs, aims to improve the semantic segmentation performance and temporal consistency by exploiting the temporal continuity of video streams.
@@
 Another family of works use an optical flow module to compute the feature correspondence between consecutive frames, and then use this flow for predictive feature learning [Ding et al.(2020)Ding, Wang, Zhou, Shi, Lu, and Luo, Zhu et al.(2017)Zhu, Xiong, Dai, Yuan, and Wei, Liu et al.(2020)Liu, Shen, Yu, and Wang, Xu et al.(2018)Xu, Fu, Yang, and Lee, Baghbaderani et al.(2024)Baghbaderani, Li, Wang, and Qi, Varghese et al.(2021)Varghese, Gujamagadi, Klingner, Kapoor, Bar, Schneider, Maag, Schlicht, Huger, and Fingscheidt, Hur and Roth(2016)].
-Our method belongs to the latter category.
-However, unlike aforementioned approaches, MCDS-VSS incorporates assumptions from the domain of moving cameras and dynamic scenes into the model design, computing interpretable intermediate representations, which lead to more temporally consistent video segmentation results.
+Our method belongs to the latter category of VSS models.
+However, unlike aforementioned approaches, MCDS-VSS incorporates assumptions from the domain of moving cameras and dynamic scenes into the model design, and computes interpretable intermediate geometry and motion-aware representations, which lead to accurate and temporally consistent video segmentation results.
 Improving Segmentation via Depth & Camera Motion Estimation:
@@
 Several works have investigated the use of depth and motion for semantic segmentation in videos.
-Approaches like [Shi et al.(2020)Shi, Lin, Wang, Hung, and Wang, Athar et al.(2023)Athar, Li, Casas, and Urtasun, Cao and Behnke(2024), Zhu et al.(2023)Zhu, Han, Cai, Borse, Ghaffari, and Porikli] segment dynamic scenes by jointly processing video frames with depth information captured by LiDAR scanners; whereas methods like [Ma et al.(2017)Ma, Stückler, Kerl, and Cremers, Kundu et al.(2020)Kundu, Yin, Fathi, Ross, Brewington, Funkhouser, and Pantofaru, Chen et al.(2021)Chen, Lin, Wang, Yang, and Cheng, Rosu et al.(2020)Rosu, Quenzel, and Behnke] use depth and pose information to enforce consistency between segmentation maps from multiple viewpoints.
-Recently, depth-aware panoptic segmentation (PS) models [Qiao et al.(2021)Qiao, Zhu, Adam, Yuille, and Chen, Petrovai and Nedevschi(2023), Yuan et al.(2022)Yuan, Li, Yang, Cheng, Zhang, Tong, Zhang, and Tao] aim to jointly solve the tasks of PS and depth estimation by extending a segmentation model with a depth decoder and conditioning its prediction using instance-masks.
-[…]
-[…]
+Approaches like [Shi et al.(2020)Shi, Lin, Wang, Hung, and Wang, Athar et al.(2023)Athar, Li, Casas, and Urtasun, Cao and Behnke(2024), Zhu et al.(2023)Zhu, Han, Cai, Borse, Ghaffari, and Porikli] segment dynamic scenes by jointly processing video frames with depth information captured by LiDAR scanners; whereas methods like [Ma et al.(2017)Ma, Stückler, Kerl, and Cremers, Kundu et al.(2020)Kundu, Yin, Fathi, Ross, Brewington, Funkhouser, and Pantofaru, Chen et al.(2021)Chen, Lin, Wang, Yang, and Cheng, Rosu et al.(2020)Rosu, Quenzel, and Behnke] use depth and camera pose information in combination with a semantic segmentation model in order to improve the segmentation performance by enforcing consistency between predictions from multiple viewpoints.
+Recently, depth-aware panoptic segmentation models [Qiao et al.(2021)Qiao, Zhu, Adam, Yuille, and Chen, Petrovai and Nedevschi(2023), Yuan et al.(2022)Yuan, Li, Yang, Cheng, Zhang, Tong, Zhang, and Tao] aim to jointly solve the tasks of panoptic segmentation and depth estimation by extending a segmentation model with a depth decoder and conditioning its prediction using instance-masks.
 […]
@@
 3 MCDS-VSS Structured Filtering Method
-We propose MCDS-VSS, illustrated in Fig.
-1, a structured filter that improves the temporal consistency of a semantic segmentation model on moving camera dynamic scene scenarios.
+We propose MCDS-VSS, illustrated in Figure 1, a structured filter that improves the temporal consistency of a semantic segmentation model on moving camera dynamic scene scenarios.
 MCDS-VSS learns in a self-supervised way to estimate geometry and motion representations, i.e., scene depth and camera ego-motion.
@@
 MCDS-VSS learns in a self-supervised manner to estimate scene geometry and camera motion, which are then used to improve the temporal consistency of a segmentation model.
-Fig.
-2 illustrates our two-step self-supervised approach for learning the scene geometry with camera motion and for distillation of object dynamics.
+Figure 2 illustrates our two-step self-supervised approach for learning the scene geometry with camera motion and for distillation of object dynamics.
 Scene Geometry and Ego-Motion: We train our model to predict the monocular depth and camera pose transformation of the vehicle in a self-supervised manner by solving a novel view-synthesis pretext task in which a target image $\mathbf{x}_{t}$ is rendered from a source image $\mathbf{x}_{t-1}$ by modeling the static scene features that change due to the ego-motion [Zhou et al.(2017)Zhou, Brown, Snavely, and Lowe, Godard et al.(2019)Godard, Mac Aodha, Firman, and Brostow].
 To predict the scene geometry, MCDS-VSS incorporates a depth decoder $\mathcal{D}_{\textrm{d}}$ that outputs the depth $\mathbf{d}_{t}$ and inverse depth $\mathbf{d}^{\triangleleft}_{t}$111 With slight abuse of notation, we denote the $\mathbf{d}^{\triangleleft}_{t}$ as disparities.
-of the scene given the input feature maps $\mathbf{h}_{t}$; whereas to compute the camera motion between two images we employ a motion encoder $\mathcal{E}_{\textrm{m}}$ that computes motion features between two sets of feature maps, and an ego-motion decoder $\mathcal{D}_{\textrm{c}}$, which predicts the camera transformation between two time steps $\textrm{C}_{{t-1}}^{\,{t}}$, parameterized as a 6-dim.
-vector containing the translation and Euler angles of the camera transformation matrix.
-We then render the ego-warped image $\mathbf{\hat{x}}^{\textrm{ego}}_{t}$ using the estimated depth and ego-motion:
+of the scene given the input feature maps $\mathbf{h}_{t}$; whereas to compute the camera motion between two images we employ a motion encoder $\mathcal{E}_{\textrm{m}}$ that computes motion features between two sets of feature maps, and an ego-motion decoder $\mathcal{D}_{\textrm{c}}$, which predicts the camera transformation between two time steps $\textrm{C}_{{t-1}}^{\,{t}}$, parameterized as a 6-dimensional vector containing the translation parameters and Euler angles of the camera transformation matrix.
+We then render the ego-warped image $\mathbf{\hat{x}}^{\textrm{ego}}_{t}$ using the estimated scene depth and ego-motion:
 $\mathbf{\hat{x}}^{\textrm{ego}}_{t}=\mathcal{F}_{\textrm{fwd}}(\mathbf{x}_{t-1},\mathbf{d}_{t-1},\textrm{C}_{{t-1}}^{\,{t}},K),$ (1)
@@
 If the depth and ego-motion estimates are accurate, the resulting warped images should match the target images except for occluded regions and moving objects.
-Therefore, as illustrated in Fig.
-2(a), we train the modules for self-supervised depth and ego-motion estimation by optimizing the following loss function:
+Therefore, as illustrated in Figure 2(a), we train the modules for self-supervised depth and ego-motion estimation by optimizing the following loss function:
 $\mathcal{L}_{\textrm{Depth}}=\mathcal{L}_{\textrm{Photo}}({\mathbf{\hat{x}}^{\textrm{ego}}_{t}},{\mathbf{x}_{t}})+\lambda_{\textrm{Reg}}\cdot\mathcal{L}_{\textrm{Reg}}({\mathbf{d}^{\triangleleft}_{t-1}}),$ (2)
@@
 To mitigate the effect of disocclusions and moving objects during training, we use the auto-masking and per-pixel minimum processing steps proposed in [Godard et al.(2019)Godard, Mac Aodha, Firman, and Brostow].
-[…]
-[…]
 (a) SSL of depth and ego-motion.
-[…]
-[…]
 (b) Distillation of object-motion.
 Figure 2: Learning geometry and motion.
-a) We learn the depth $\mathbf{d}_{t-1}$ and ego-motion $\textrm{C}_{{t-1}}^{\,{t}}$ given two video frames by enforcing a photometric loss between warped and target frame $\mathcal{L}_{\textrm{Photo}}$ and a depth regularization $\mathcal{L}_{\textrm{Reg}}$.
-b) Given an ego-warped image, we train a flow decoder to model the dynamics of moving objects by distilling a pretrained RAFT model.
+a) We learn the scene depth $\mathbf{d}_{t-1}$ and ego-motion $\textrm{C}_{{t-1}}^{\,{t}}$ in a self-supervised manner given two video frames by enforcing a photometric loss $\mathcal{L}_{\textrm{Photo}}$ between the ego-warped $\mathbf{\hat{x}}^{\textrm{ego}}_{t}$ and target frames $\mathbf{x}_{t}$, as well as a depth regularization $\mathcal{L}_{\textrm{Reg}}$.
+b) Given an ego-warped image, we train a residual flow decoder to predict the residual optical flow $\hat{\textrm{F}}_{{t-1}}^{\,{t}}$ that parameterizes the dynamics of moving objects in the scene by distilling a pretrained RAFT model.
 Object Motion: Assuming static scenes as well as accurate depth and ego-motion estimates, the predicted ego-warped images $\mathbf{\hat{x}}^{\textrm{ego}}_{t}$ are identical, up to occluded regions, to the target images $\mathbf{x}_{t}$.
@@
 driving cars or pedestrians).
-As illustrated in Fig.
-2(b), we estimate the residual optical flow $\hat{\textrm{F}}_{{t-1}}^{\,{t}}$ between the ego-warped and target images, which parameterizes the dynamics of moving objects, by training an residual flow decoder $\mathcal{R}_{\textrm{f}}$ while keeping all other modules frozen.
+As illustrated in Figure 2(b), we estimate the residual optical flow $\hat{\textrm{F}}_{{t-1}}^{\,{t}}$ between the ego-warped and target images, which encodes the dynamics of moving objects, by training a residual flow decoder $\mathcal{R}_{\textrm{f}}$ while keeping all other modules frozen.
+The residual flow $\hat{\textrm{F}}_{{t-1}}^{\,{t}}$ is parameterized as a 2D flow field that encodes the per-pixel motion in the horizontal and vertical directions needed to align the ego-warped images $\mathbf{\hat{x}}^{\textrm{ego}}_{t}$ to the corresponding target images $\mathbf{x}_{t}$.
 $\mathcal{R}_{\textrm{f}}$ is trained to match the optical flow predictions of the large state-of-the-art optical flow model RAFT [Teed and Deng(2020)]:
@@
 4 subsec: ssl
-form the core of the MCDS-VSS structured filter, which is depicted in Fig.
-1.
-It propagates information over time using two different filter states, namely a scene state $\mathbf{s}$ that encodes the scene contents and geometry and a camera state $\mathbf{c}$, which encodes the ego-motion of the vehicle.
+form the core of the MCDS-VSS structured filter, which is depicted in Figure 1.
+It propagates information over time using two different filter states, namely a scene state $\mathbf{s}$ that encodes the scene contents and geometry, and a camera state $\mathbf{c}$ that encodes the ego-motion of the vehicle.
 It consists of six components: ego-motion filter, depth estimation, ego-motion compensation, residual flow estimation, object motion compensation, and feature fusion.
-Ego-Motion Filter: The ego-motion filter extends the motion encoder and camera predictor modules in order to aggregate motion information over time and enforce the prediction of temporally consistent ego-motion.
-The temporal integration is achieved via a motion update module, which is implemented as a ConvGRU [Ballas et al.(2016)Ballas, Yao, Pal, and Courville] recurrent layer that jointly processes the motion features and previous camera state $\mathbf{c}_{t-1}$ and outputs an updated camera state $\mathbf{c}_{t}$ from which the ego-motion can be then predicted:
+Ego-Motion Filter: The ego-motion filter extends the motion encoder $\mathcal{E}_{\textrm{m}}$ and ego-motion decoder $\mathcal{D}_{\textrm{c}}$ modules in order to aggregate motion information over time and enforce the prediction of temporally consistent ego-motion.
+The temporal integration is achieved via a motion update module, which is implemented as a ConvGRU [Ballas et al.(2016)Ballas, Yao, Pal, and Courville] recurrent layer that jointly processes the motion features and previous camera state $\mathbf{c}_{t-1}$, and outputs an updated camera state $\mathbf{c}_{t}$ from which the ego-motion can be then predicted:
 $\mathbf{c}_{t}=\text{ConvGRU}(\mathcal{E}_{\textrm{m}}(\mathbf{h}_{t},\mathbf{h}_{t-1}),\mathbf{c}_{t-1}),\hskip 28.45274pt\textrm{C}_{{t-1}}^{\,{t}}=\mathcal{D}_{\textrm{c}}(\mathbf{c}_{t}).$ (6)
-Depth Estimation and Ego-Motion Compensation: Given the past scene state $\mathbf{s}_{t-1}$, the depth decoder $\mathcal{D}_{\textrm{d}}$ computes the depth map $\mathbf{d}_{t-1}$.
-This scene geometry and the ego-motion $\textrm{C}_{{t-1}}^{\,{t}}$ are used as in Eq.
-(1) to project $\mathbf{s}_{t-1}$ to time $t$.
+Depth Estimation and Ego-Motion Compensation:.
+Given the past scene state $\mathbf{s}_{t-1}$, the depth decoder $\mathcal{D}_{\textrm{d}}$ computes the depth map $\mathbf{d}_{t-1}$.
+This scene geometry and the estimated ego-motion $\textrm{C}_{{t-1}}^{\,{t}}$ are used as in Equation (1) to project $\mathbf{s}_{t-1}$ to time $t$.
 The resulting ego-warped scene state $\mathbf{s}^{\textrm{ego}}_{t}$ encodes scene contents and geometry after compensation for ego-motion.
-Object-Motion Compensation: This module models dynamic objects in the scene, such as pedestrians or vehicles, and updates the scene state to compensate for the motion of such objects.
+Residual Flow Estimation and Object Motion Compensation: These modules model dynamic objects in the scene, such as pedestrians or vehicles, and update the scene state to compensate for the motion of such objects.
 We jointly process the ego-warped scene state $\mathbf{s}^{\textrm{ego}}_{t}$ and the current image features $\mathbf{h}_{t}$ with the residual flow decoder $\mathcal{R}_{\textrm{f}}$ in order to compute the residual flow $\hat{\textrm{F}}_{{t-1}}^{\,{t}}$, which represents the pixel displacement of moving objects between consecutive time steps.
@@
 depth or object flow).
-To ease the training process, we propose a multi-stage approach in which we first train the encoder and decoder modules using image pairs or triplets, and then integrate and train the filter modules using sequences of six frames in order to improve segmentation performance and temporal consistency while retaining interpretable representations.
-Table 1: Training stages and hyper-parameters.
+To ease the training process, we propose a multi-stage training procedure in which we first train the encoder and decoder modules using image pairs or triplets, and then integrate and train the filter modules using sequences of six frames in order to gather scene context information and improve the segmentation performance and temporal consistency while retaining interpretable representations.
+Table 1: MCDS-VSS training stages and hyper-parameters.
 Stage Training Goal Loss Function LR # Imgs
-1 Segmentation & SSL Geometry $\mathcal{L}_{\textrm{Seg}}+\lambda_{\textrm{D}}\cdot\mathcal{L}_{\textrm{Depth}}$   (2) $2\cdot 10^{-4}$ 3
+1 Segmentation & SSL Geometry $\mathcal{L}_{\textrm{Segm}}+\lambda_{\textrm{D}}\cdot\mathcal{L}_{\textrm{Depth}}$   (2) $2\cdot 10^{-4}$ 3
 2 Distillation of Object Motion $\mathcal{L}_{\textrm{Flow}}$   (5) $1\cdot 10^{-4}$ 2
 3 Ego-Motion Filter $\mathcal{L}_{\textrm{Ego}}$   (9) $8\cdot 10^{-5}$ 6
-4 Temporal Integration $\mathcal{L}_{\textrm{Seg}}+\lambda_{\textrm{TC}}\cdot\mathcal{L}_{\textrm{TC}}$   (10) $8\cdot 10^{-5}$ 6
+4 Temporal Integration $\mathcal{L}_{\textrm{Segm}}+\lambda_{\textrm{TC}}\cdot\mathcal{L}_{\textrm{TC}}$   (10) $8\cdot 10^{-5}$ 6
 MCDS-VSS undergoes a four-stage training process, outlined in Table 1. Initially, as detailed in
 5 subsec: ssl
-, MCDS-VSS encoder and decoder modules are jointly trained for self-supervised learning of geometry and ego-motion, as well as for semantic segmentation by minimizing a combination of cross entropy $\mathcal{L}_{\textrm{Seg}}$ and SSL geometry $\mathcal{L}_{\textrm{Depth}}$ losses.
-Following [Godard et al.(2019)Godard, Mac Aodha, Firman, and Brostow], we use image triplets $(\mathbf{x}_{t-\tau},\mathbf{x}_{t},\mathbf{x}_{t+\tau})$, with $\mathbf{x}_{t}$ being the target.
-Subsequently, we train the residual flow decoder as described in
+, MCDS-VSS encoder and decoder modules are jointly trained for self-supervised learning of geometry and ego-motion, as well as for semantic segmentation by minimizing a combination of cross entropy $\mathcal{L}_{\textrm{Segm}}$ and SSL geometry $\mathcal{L}_{\textrm{Depth}}$ losses.
+Following [Godard et al.(2019)Godard, Mac Aodha, Firman, and Brostow], we use image triplets $(\mathbf{x}_{t-\tau},\mathbf{x}_{t},\mathbf{x}_{t+\tau})$, with $\mathbf{x}_{t}$ being the target image and $\tau$ being the temporal distance between source and target frames used during this first training stage.
+Subsequently, we train the residual flow decoder using image pairs as described in
 6 subsec: ssl
 while keeping the remaining modules frozen.
-In the third stage, with the goal of improving the temporal continuity of the predicted ego-motion, we train the ego-motion filter and compensation by minimizing the loss function:
+In the third stage, with the goal of improving the temporal continuity of the predicted ego-motion, we train the ego-motion filter and compensation using short video sequences of length $T$ by minimizing the loss function:
 $\mathcal{L}_{\textrm{Ego}}=\frac{1}{T}\sum_{t=1}^{T}\mathcal{L}_{\textrm{Photo}}({\mathbf{\hat{x}}^{\textrm{ego}}_{t}},{\mathbf{x}_{t}}),$ (9)
 which enforces the model to compute accurate camera motion estimates in order to align the ego-warped state with the current observations.
-Finally, we jointly train the fusion module and fine-tune the segmentation decoder by minimizing the following loss function:
-$\mathcal{L}=\frac{1}{T}\sum_{t=1}^{T}\mathcal{L}_{\textrm{Seg}}({\mathbf{\hat{y}}_{t}},{\mathbf{y}_{t}})+\lambda_{\textrm{TC}}\cdot\mathcal{L}_{\textrm{TC}}({\mathbf{\tilde{y}}_{t}},{\mathbf{\hat{y}}_{t}}),$ (10)
-where $\mathcal{L}_{\textrm{Seg}}$ is the cross entropy loss function and $\mathcal{L}_{\textrm{TC}}$ is a temporal consistency regularizer that enforces the segmentation $\mathbf{\tilde{y}}_{t}$ computed by decoding $\mathbf{s}^{\textrm{full}}_{t}$ to be close to the actual predicted segmentation maps $\mathbf{\hat{y}}_{t}$.
+Finally, in the last training stage we jointly train the feature fusion module and fine-tune the segmentation decoder by minimizing the following loss function:
+$\mathcal{L}=\frac{1}{T}\sum_{t=1}^{T}\mathcal{L}_{\textrm{Segm}}({\mathbf{\hat{y}}_{t}},{\mathbf{y}_{t}})+\lambda_{\textrm{TC}}\cdot\mathcal{L}_{\textrm{TC}}({\mathbf{\tilde{y}}_{t}},{\mathbf{\hat{y}}_{t}}),$ (10)
+where $\mathcal{L}_{\textrm{Segm}}$ is the cross entropy loss function and $\mathcal{L}_{\textrm{TC}}$ is a temporal consistency regularizer that enforces the segmentation $\mathbf{\tilde{y}}_{t}$ computed by decoding $\mathbf{s}^{\textrm{full}}_{t}$ to be close to the actual predicted segmentation maps $\mathbf{\hat{y}}_{t}$.
 7 Experimental Evaluation
 7.1 Experiment Setup
-Dataset: We evaluate MCDS-VSS on the Cityscapes [Cordts et al.(2016)Cordts, Omran, Ramos, Rehfeld, Enzweiler, Benenson, Franke, Roth, and Schiele] dataset, which contains 5,000 automotive video sequences, each containing 30 images of size 1024$\times$2048, recorded in 50 German cities, where only the 20th frame is annotated.
+Dataset: We evaluate MCDS-VSS on the Cityscapes [Cordts et al.(2016)Cordts, Omran, Ramos, Rehfeld, Enzweiler, Benenson, Franke, Roth, and Schiele] dataset, which contains 5,000 automotive video sequences recorded in 50 German cities.
+Each sequence contains 30 images of size 1024$\times$2048, where only the 20th frame is annotated.
 This dataset is a good benchmark for our model, since it contains real-world dynamic scenes recorded from a moving vehicle.
-We augment the data using color jittering, mirroring, and random cropping.
+We augment the data using color jittering, mirroring and random cropping.
 Evaluation Metrics: We evaluate the segmentation performance and temporal consistency of our model.
@@
 Further implementation details are provided in Appendix B.
-7.2 Comparison with Existing Methods
-In Table 2, we quantitatively compare MCDS-VSS with several existing image and video segmentation models using small (left) and larger (right) backbones.
-For both variants, MCDS-VSS achieves the highest temporal consistency among all compared methods, while retaining a competitive segmentation performance.
-Furthermore, in contrast to other approaches aiming to improve the TC of a segmentation model, e.g.
-TCNet [Varghese et al.(2021)Varghese, Gujamagadi, Klingner, Kapoor, Bar, Schneider, Maag, Schlicht, Huger, and Fingscheidt], MCDS-VSS does not sacrifice segmentation performance in order to improve the temporal consistency, outperforming multiple VSS models for both backbone variants.
-Table 2: Comparison of segmentation models on the Cityscapes validation set using different backbones. Best two results are highlighted in boldface and underlined, respectively.
+Table 2: Comparison of image and video segmentation models on the Cityscapes validation set using small (left) and larger (right) backbones.
+We evaluate the segmentation accuracy (mIoU) and temporal consistency (TC) of the models.
+Best two results are highlighted in boldface and underlined, respectively.
 Cityscapes
@@
 MCDS-VSS (ours) ResNet18 75.1 74.5
-[…]
-[…]
-Figure 3: Qualitative evaluation.
+7.2 Comparison with Existing Methods
+In Table 2, we quantitatively compare MCDS-VSS with several existing image and video segmentation models using small (left) and larger (right) backbones.
+For both variants, MCDS-VSS achieves the highest temporal consistency among all compared methods, while retaining a competitive segmentation performance.
+Furthermore, in contrast to other approaches aiming to improve the TC of a segmentation model, e.g.
+TCNet [Varghese et al.(2021)Varghese, Gujamagadi, Klingner, Kapoor, Bar, Schneider, Maag, Schlicht, Huger, and Fingscheidt], MCDS-VSS does not sacrifice segmentation performance in order to improve the temporal consistency, outperforming multiple VSS models for both backbone variants.
+Figure 3: Qualitative evaluation on a validation sequence of five frames.
 a) Input frames, b) HRNetV2, c) MCDS-VSS (ours), d) Estimated scene depth, e) Estimated residual flow.
-We highlight areas of the segmentation masks where MCDS-VSS obtains visibly more accurate and consistent segmentations.
-In Fig.
-3, we show a qualitative result comparing MCDS-VSS with an HRNetV2 baseline.
-Our method achieves more accurate and temporally consistent segmentations compared to HRNet, correctly segmenting the traffic signs and reducing the amount of flickering between frames.
-Furthermore, we show MCDS-VSS interpretable intermediate representations, such as the depth of the scene and the residual flow, which encodes the movement of the pedestrians in the scene, as well as some corrections for other objects, such as the hood of the ego-vehicle.
+We highlight areas of the segmentation masks where MCDS-VSS obtains visibly more accurate and temporally consistent segmentations, such as the traffic signs or the bus, which HRNetV2 mislabels as truck.
+In Figure 3, we show a qualitative result comparing MCDS-VSS with the HRNetV2 baseline on a validation sequence of five frames.
+Whereas the baseline mislabels the bus as a truck and outputs inconsistent segmentation labels on certain regions such as the traffic signs, our method achieves more accurate and temporally consistent segmentations, predicting more stable semantic labels across video frames.
+Furthermore, we show MCDS-VSS interpretable intermediate representations, such as the estimated scene depth and residual optical flow, which encodes the movement of the vehicles in the scene, as well as corrections for the hood of the ego-vehicle.
 Further visualizations can be found in Appendix D.
@@
 , we compare MCDS-VSS with different filter designs, including unstructured RNNs (ConvGRU) [Siam et al.(2017)Siam, Valipour, Jagersand, and Ray, Pfeuffer et al.(2019)Pfeuffer, Schulz, and Dietmayer], optical-flow based filters [Gadde et al.(2017)Gadde, Jampani, and Gehler] (flow-only), and a MCDS-VSS variant modeling only the static scene features (geom-only) [Wagner et al.(2018)Wagner, Fischer, Herman, and Behnke].
-The results, reported in Table 3, show that MCDS-VSS, which decouples the modeling of static and dynamic scene features, achieves the best segmentation performance and temporal consistency.
-Model Ablation: In Table 4 we measure the effect that our SSL training and MCDS-VSS filter have on the segmentation performance and temporal consistency.
+The results, reported in Table 3, show that filter designs that project scene features using geometry and motion representations outperform the ConvGRU, which learns to model the video dynamics solely from data.
+Furthermore, MCDS-VSS, which decouples the modeling of static and dynamic scene features, achieves the best segmentation performance and temporal consistency among the compared filter designs.
+Model Ablation: In Table 4 we measure the effect that our joint training procedure of semantic segmentation and SSL depth and ego-motion, as well as the MCDS-VSS filter have on the segmentation performance and temporal consistency.
 For two different segmentation models, i.e.
-DeepLabV3+ and HRNetV2, we compare the results after each training stage with those of the model trained for image segmentation only.
-First, we note that SSL training improves the temporal consistency without significantly compromising the segmentation performance.
-We argue that jointly learning SSL depth and motion with semantic segmentation allows the model to encode the input frames into more robust geometry-aware representations.
+DeepLabV3+ with a ResNet18 backbone and HRNetV2, we compare the results after each training stage with those of the model trained for image segmentation only.
+First, we note that jointly learning semantic segmentation with SSL depth and ego-motion estimation improves the temporal consistency without significantly compromising the segmentation performance.
+We argue that the joint training procedure allows the model to encode the input frames into more robust geometry-aware representations.
 Finally, the MCDS-VSS filter significantly improves the temporal consistency (>4.6% w.r.t base model), while almost matching the segmentation performance of the base DeepLabV3+ model, and even outperforming HRNetV2.
@@
 It also estimates the motion of external objects and leverages these representations to improve the temporal consistency of a semantic segmentation model without sacrificing segmentation performance.
-MCDS-VSS follows a prediction-fusion approach in which scene geometry and camera motion are first used to compensate for ego-motion, then residual flow is used to compensate the motion of dynamic objects, and finally the projected features are fused with the current observations.
-In our experiments, we showed that MCDS-VSS outperforms multiple VSS baselines on Cityscapes—achieving superior temporal consistency and parsing the scene into human-interpretable representations.
-Acknowledgments
+MCDS-VSS follows a prediction-fusion approach in which scene geometry and camera motion are first used to compensate for ego-motion, then residual flow is used to compensate the motion of dynamic objects, and finally the projected features are fused with the current observations in order to obtain a temporally consistent representation of the scene.
+In our experiments, we showed that MCDS-VSS outperforms multiple VSS baselines on Cityscapes—achieving superior segmentation temporal consistency and parsing the scene into human-interpretable representations, such as depth, ego-motion and object flow.
+Acknowledgment
 This work was funded by grant BE 2556/16-2 (Research Unit FOR 2535 Anticipating Human Behavior) of the German Research Foundation (DFG).
@@
 To seamlessly integrate this module into our MCDS-VSS filter, we modify the implementation of raft_small provided by PyTorch [^2] https://pytorch.org/vision/main/models/raft.html by replacing the expensive feature and context encoders with a single convolutional block that directly processes the ego-warped $\mathbf{s}^{\textrm{ego}}_{t}$ and image features $\mathbf{h}_{t}$.
-We set the number of refinement iterations to 10.
+We set the number of refinement iterations to 12.
 MCDS-VSS Filter: We instantiate the MCDS-VSS filter using the modules described above, after being pretrained for semantic segmentation, SSL of depth and ego-motion, and distillation of object motion.
@@
 We experimented with learning the initial state representations; however it did not yield any qualitative or quantitative improvements, while increasing the number of learnable parameters.
-Table 7: Training stages and hyper-parameters.
-Stage Training Goal Loss Function LR Batch Train Time # Imgs
-1 Segmentation & SSL Geometry $\mathcal{L}_{\textrm{Seg}}+\lambda_{\textrm{D}}\cdot\mathcal{L}_{\textrm{Depth}}$ $2\cdot 10^{-4}$ 12 40h 3
-2 Distillation of Object Motion $\mathcal{L}_{\textrm{Flow}}$ $1\cdot 10^{-4}$ 4 18h 2
-3 Ego-Motion Filter $\mathcal{L}_{\textrm{Ego}}$ $8\cdot 10^{-5}$ 8 8h 6
-4 Temporal Integration $\mathcal{L}_{\textrm{Seg}}+\lambda_{\textrm{TC}}\cdot\mathcal{L}_{\textrm{TC}}$ $8\cdot 10^{-5}$ 4 12h 6
 B.2 Training and Inference
@@
 We empirically set the loss weight values to $\lambda_{\textrm{D}}=2$, $\lambda_{\textrm{Reg}}=10^{-5}$ and $\lambda_{\textrm{TC}}=1$.
-Tables 8 and 9 report the number of learnable parameters, throughput and inference time for each individual module, as well as for the complete model, for the MCDS-VSS variants based on DeepLabV3+ and HRNetV2, respectively.
-We emphasize that the ego-motion and object motion compensation modules do not have any learnable parameters, but instead hard-wire our knowledge from the moving camera dynamic scene domain to project the previous scene features into the current time-step using geometry and motion representations.
-We also observe that MCDS-VSS inference is severely limited by its residual flow decoder.
-Adapting such module to exploit recent advances in fast optical flow estimation [Kong et al.(2021)Kong, Shen, and Yang]), as well as using more efficient image encoders [Howard et al.(2019)Howard, Sandler, Chu, Chen, Chen, Tan, Wang, Zhu, Pang, Vasudevan, et al.], could allow MCDS-VSS to be used for real time video semantic segmentation.
+Table 7: MCDS-VSS training stages and hyper-parameters.
+Stage Training Goal Loss Function LR Batch Train Time # Imgs
+1 Segmentation & SSL Geometry $\mathcal{L}_{\textrm{Segm}}+\lambda_{\textrm{D}}\cdot\mathcal{L}_{\textrm{Depth}}$ $2\cdot 10^{-4}$ 12 40h 3
+2 Distillation of Object Motion $\mathcal{L}_{\textrm{Flow}}$ $1\cdot 10^{-4}$ 4 18h 2
+3 Ego-Motion Filter $\mathcal{L}_{\textrm{Ego}}$ $8\cdot 10^{-5}$ 8 8h 6
+4 Temporal Integration $\mathcal{L}_{\textrm{Segm}}+\lambda_{\textrm{TC}}\cdot\mathcal{L}_{\textrm{TC}}$ $8\cdot 10^{-5}$ 4 12h 6
 Table 8: Throughput, inference speed (in ms) and number of learnable parameters for MCDS-VSS based on DeepLabV3+.
@@
 Total MCDS-VSS 26.2M 5.6 177.6
-Appendix C Additional Quantitative Results
+Tables 8 and 9 report the number of learnable parameters, throughput and inference time for each individual module, as well as for the complete model, for the MCDS-VSS variants based on DeepLabV3+ and HRNetV2, respectively.
+We emphasize that the ego-motion and object motion compensation modules do not have any learnable parameters, but instead hard-wire our knowledge from the moving camera dynamic scene domain to project the previous scene features into the current time-step using geometry and motion representations.
+We also observe that MCDS-VSS inference is severely limited by its residual flow decoder.
+Adapting such module to exploit recent advances in fast optical flow estimation [Kong et al.(2021)Kong, Shen, and Yang]), as well as using more efficient image encoders [Howard et al.(2019)Howard, Sandler, Chu, Chen, Chen, Tan, Wang, Zhu, Pang, Vasudevan, et al.], could allow MCDS-VSS to be used for real time video semantic segmentation.
+Appendix C Quantitative Results
 In Table 10 we compare for individual classes of the Cityscapes dataset the segmentation performance and temporal consistency of MCDS-VSS with an HRNetV2 baseline trained for semantic segmentation and SSL of geometry and motion.
@@
 Ours 98.9 88.2 92.1 64.1 65.9 69.1 69.2 79.5 70.5 67.8 90.5 65.2 69.6 48.3 53.6 71.0 75.3
-Appendix D Additional Qualitative Results
+Appendix D Qualitative Results
 D.1 Effect of Each MCDS-VSS Stage
-In Fig.
-4 we display the semantic segmentations obtained when decoding the scene features from different stages of our MCDS-VSS filter.
-We can observe how the segmentations after ego-motion compensation (Fig.
-4 b) atone for the movement of the ego-vehicle, correctly representing static scene features such as buildings or the bicycle.
+In Figure 4 we display the semantic segmentations obtained when decoding the scene features from different stages of our MCDS-VSS filter.
+We can observe how the segmentations after ego-motion compensation (Figure 4 b) atone for the movement of the ego-vehicle, correctly representing static scene features such as buildings or the bicycle.
 However, the dynamics of moving objects (e.g.
 yellow car) are not addressed in this step, thus not compensating for such movement.
-This limitation is addressed in the object motion compensation step (Fig.
-4 c).
+This limitation is addressed in the object motion compensation step (Figure 4 c).
 However, the disocclusions resulting from the moving car and inaccuracies in the residual flow estimation can lead to segmentation errors.
-Finally, fusing the projected scene state features with the current observations (Fig.
-4 d) leads to a more accurate segmentation of the scene.
-In (Fig.
-4 e) we visualize the update gate mask $\mathbf{u}$ employed for feature fusion.
+Finally, fusing the projected scene state features with the current observations (Figure 4 d) leads to a more accurate segmentation of the scene.
+In (Figure 4 e) we visualize the update gate mask $\mathbf{u}$ employed for feature fusion.
 For visualization purposes, we take the mean across all channels and assign lighter colors to spatial locations where MCDS-VSS relies on the propagated state $\mathbf{s}^{\textrm{full}}_{t}$, whereas darker colors correspond to the current features $\mathbf{h}_{t}$.
-We observe that MCDS-VSS relies on the scene state to represent static areas such as buildings or the street, whereas it relies heavier on observations for accurately segmenting fast moving objects or thin structures (e.g.
+We observe that MCDS-VSS relies on the scene state to represent static areas such as buildings or the street, whereas it relies heavier on observations for accurately segmenting disoccluded areas of the image, fast moving objects or thin structures (e.g.
 poles or street signs).
-[…]
-[…]
 Figure 4: Video segmentation for each stage in MCDS-VSS.
 a) Input images, b) segmentation after ego-motion compensation, c) segmentation after object motion compensation, d) segmentation after feature fusion, e) feature fusion update mask, lighter colors mean that filter information is used, whereas darker ones correspond to observations.
-[…]
-[…]
+D.2 MCDS-VSS Qualitative Results
+In Figure 5, we show a qualitative result comparing MCDS-VSS with the HRNetV2 baseline.
+Our method achieves more accurate and temporally consistent segmentations compared to HRNet, correctly segmenting the traffic signs and reducing the amount of flickering between frames.
+In Figures 69, we show on four validation sequences how MCDS-VSS obtains an accurate and temporally consistent segmentation of the scene, estimates the scene depth, and computes the residual motion flow, which encodes the movement of dynamic objects in the scene, as well as some minor motion corrections for other objects and scene features.
+D.3 Cross-Dataset Evaluation
+We qualitatively evaluate the robustness of MCDS-VSS by performing a cross-dataset validation in which a DeepLabV3+ baseline and our MCDS-VSS model trained on Cityscapes are qualitatively evaluated without retraining on sequences from the KITTI [Geiger et al.(2012)Geiger, Lenz, and Urtasun] dataset.
+Figures 10 and 11 illustrate the semantic segmentation predictions of both models, as well as the MCDS-VSS depth estimates on two validation sequences of the KITTI dataset.
+Due to differences with respect to the training data in the camera model and calibration, as well as different image resolution and aspect ratio, the segmentation performance of both models on the KITTI dataset is severely degraded with respect to Cityscapes.
+However, MCDS-VSS achieves a more accurate and temporally consistent video segmentation, thus verifying that incorporating geometry and motion inductive biases from the moving camera dynamic scene domain into the VSS model design leads to more robust representations and segmentation results.
+D.4 Point Clouds
+Figures 1215 show examples of RGB and semantic point clouds rendered by backprojecting image values and semantic labels using the depth maps estimated by MCDS-VSS and known camera intrinsics.
+The high-quality depth maps computed by MCDS-VSS allow for an accurate 3D representation of the scene.
+Figure 5: Qualitative evaluation.
+a) Input frames, b) HRNetV2, c) MCDS-VSS , d) Estimated scene depth, e) Estimated residual flow.
+We highlight areas of the segmentation masks where MCDS-VSS obtains visibly more accurate and temporally consistent segmentations.
 Figure 6: MCDS-VSS qualitative evaluation. a) Input frames, b) semantic segmentation, c) scene depth, d) residual flow.
-[…]
-[…]
 Figure 7: MCDS-VSS qualitative evaluation. a) Input frames, b) semantic segmentation, c) scene depth, d) residual flow.
-[…]
-[…]
 Figure 8: MCDS-VSS qualitative evaluation. a) Input frames, b) semantic segmentation, c) scene depth, d) residual flow.
-[…]
-[…]
 Figure 9: MCDS-VSS qualitative evaluation. a) Input frames, b) semantic segmentation, c) scene depth, d) residual flow.
-D.2 MCDS-VSS Qualitative Results
-Despite being trained without any depth or motion labels, and solely with semantic segmentation annotations for a single frame per video, our model parses the scene into semantic, geometry and motion representations.
-In Figs.
-58, we show on four validation sequences how MCDS-VSS obtains an accurate and temporally consistent segmentation of the scene, estimates the scene depth, and computes the residual motion flow, which encodes the movement of dynamic objects in the scene, as well as some minor motion corrections for other objects and scene features.
-D.3 Point Clouds
-Figs.
-912 show examples of RGB and semantic point clouds rendered by backprojecting image values and semantic labels using the depth maps estimated by MCDS-VSS and known camera intrinsics.
-The high-quality depth maps computed by MCDS-VSS allow for an accurate 3D representation of the scene.
-[…]
-[…]
+Figure 10: Cross-dataset qualitative evaluation of models trained on Cityscapes and evaluated on KITTI. a) Input frames, b) DeepLabV3+ baseline, c) MCDS-VSS (ours), d) estimated scene depth.
+Figure 11: Cross-dataset qualitative evaluation of models trained on Cityscapes and evaluated on KITTI. a) Input frames, b) DeepLabV3+ baseline, c) MCDS-VSS (ours), d) estimated scene depth.
 Figure 12: RGB and semantic point clouds rendered by lifting image values and semantic labels to 3D space.
-[…]
-[…]
 Figure 13: RGB and semantic point clouds rendered by lifting image values and semantic labels to 3D space.
-[…]
-[…]
 Figure 14: RGB and semantic point clouds rendered by lifting image values and semantic labels to 3D space.
-[…]
-[…]
 Figure 15: RGB and semantic point clouds rendered by lifting image values and semantic labels to 3D space.